(1) Evaluate the following limit using lHospitals rule. (Hint: Implicit differentiation and properties of logarithms) lim
Answer #1

Given:-\ \lim x \to \infty\ \big(4+\frac{1}{x}\big)^{\frac{-2}{x}}\\ Let\ y=\lim x \to \infty\ \big(4+\frac{1}{x}\big)^{\frac{-2}{x}}\\ Apply\ \log\ we\ have\\ ln(y) = \lim x \to \infty\ ln\big(4+\frac{1}{x}\big)^{\frac{-2}{x}}\\ =\lim x \to \infty\ \frac{-2}{x} ln\big(4+\frac{1}{x}\big)\\ Apply\ L'hospital\ we\ have\\ =\lim x \to \infty\ \frac{-2}{\big(4+\frac{1}{x}\big)}\big(\frac{-1}{x^{2}}\big)\\ ln (y) =0\ as\ x \to \infty\\ y= e^{0}=1\ as\ x \to\infty\\ So\ we\ have \lim x \to \infty\ \big(4+\frac{1}{x}\big)^{\frac{-2}{x}}=1